State Green s theorem and verify for ∮c(x²+y²)dx+(3y+2x)dy ,where C is triangle with vertices at (0,0),(1,0),(1,1)


1
Expert's answer
2021-08-12T07:27:28-0400

According to Green's theorem, the line integral equals the double integral of this quantity across the contained region. We might begin by assuming that the region is both vertically and horizontally simple. As a result, the two line integrals across this line cancel each other out.

"∮_c(x²+y²)dx+(3y+2x)dy\\
\int_{x=0}^1\int_{y=0}^1\{ \frac{\partial}{\partial x}(3y+2x)- \frac{\partial}{\partial y}(x^2+y^2)\}dydx\\
 =\int _0^1\int _0^12-\frac{\partial \:}{\partial \:y}\left(x^2+y^2\right)dydx\\
=\int _0^1\int _0^12-2ydydx\\
=\int _0^11dx\\
=1"